OpenStudy (anonymous):

3 2/3 + 4 3/4? please explain

OpenStudy (nikato):

u can ignore the 3 and 4 for now. so your just looking at 2/3 + 3/4 because they have different denominators, the first thing to do is to somehow make them have common denominators, so you can add them up. do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

add 3 and 4 and find the common denominator between 3 and 4 would be 12 so it becomes 8/12+9/12= 17/12 or 1 and 5/12

OpenStudy (anonymous):

so the final answer is 8 and5/12

OpenStudy (timaashorty):

Add the whole numbers: 3+4 = 7 Now Change the fractions with the same denominator (least common multiple) since they both have different donominators. I put 12 since both 3 and 4 multiply 12. 2/3 = 6/12 3/4= 9/12 now add 6/12 and 9/12 6+9= 15/12 7 15 / 12 = \[8 \frac{ 5 }{ 12 }\]

OpenStudy (anonymous):

thank you guys so so much!!!!

OpenStudy (timaashorty):

you're welcome [:

OpenStudy (anonymous):

no problem and remember when adding subtracting or multiplying fractions you find a common denominator or least common denominator and it makes it loads easier and make sure its in simplest form or you will get marked down i hope this helps @Memyselfmaria and if you ever need my help just message me or mention me under your question by typing @torchriswil

OpenStudy (anonymous):

sure thing! thank you again!

OpenStudy (anonymous):

no problem happy to help

OpenStudy (anonymous):

\[3\frac{ 2 }{3 }=\frac{ 3*3+2 }{3 }=\frac{ 11 }{ 3 }\] \[4\frac{ 3 }{ 4 }=\frac{ 4*4+3 }{ 4 }=\frac{ 19 }{4 }\] L.C.M.of 3 ND 4=12 \[\frac{ 11 }{3 }+\frac{ 19 }{4 }=\frac{ 11*4+19*3 }{12 }=\frac{ 44+57 }{12 }=\frac{ 101 }{12 }=8\frac{ 5 }{12 }\]